• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 05:26:23 on April18,2004

    Entry number 118448

    Followups:

    keyword=target alarms

    the target alarms just went off. it's complaining about stuff in loops 1 and 2. the alarms were temporary except for L2_CT_loop_R. this thing keeps creeping over 85 and then comes back down.

    the loop 3 target heater went red. now it is okay. i didn't see a spike in the loop 3 heater stipchart. i didn't see a spike in the 15K coolant or the ESR hall A 15K supply temp either. everything looks normal in loop 3. in addition it is out the the beam so it should be stable (wrt beam trips).

    i will keep an eye on the loop 2 temp warnings...


    FIGURE 1

    FIGURE 2